# Characterizations of amorphic schemes and fusions of pairs

### v2

@@ 2.3 Strongly regular graphs
We now mention a few known facts. First of all, the complement of a strongly regular graph of a (negative) Latin square type is a strongly regular graph of the same type (which is easily checked using the eigenvalues). Secondly, the union of two edge-disjoint and commuting strongly regular graphs of the same type is again a strongly regular graph of the same type. Again, this follows from the eigenvalues and the second part of the following observation, which follows from Lemma 2.2.
@@ 4 Schemes with many (negative) Latin square type relations
Proof 5
We note that if all but possibly one relation is of the same type, then the remaining relation, which is the complement of the union of all others must also be of the same type, and hence the scheme is amorphic [14]. We may thus assume that there is at least one relation of strictly Latin square type, say $A_{1}$, and at least one relation of strictly negative Latin square type, say $A_{2}$.
@@ 7.1 Strongly regular idempotents
Secondly, $vE_{1}\circ E_{1}=q^{0}_{11}E_{0}+q^{1}_{11}E_{1}+q^{2}_{11}E_{2}=q^{2}_{11}I+\frac{1}{v}(q^{0}_{11}-q^{2}_{11})J+(q^{1}_{11}-q^{2}_{11})E_{1}$. By multiplying this equation entrywise, once by $A_{1}$ and once by $A_{2}$, we obtain that $x^{2}-(q^{1}_{11}-q^{2}_{11})x-(q^{0}_{11}-q^{2}_{11})=0$ has solutions for $x=a,b$. Thus, $ab=q^{2}_{11}-q^{0}_{11}=q^{2}_{11}-m$ and $a+b=q^{1}_{11}-q^{2}_{11}$.

### v3

@@ 2.3 Strongly regular graphs
The following follows from Lemma 2.2.
@@
For completeness we also mention the following well-known facts.
Lemma 2.4
  • 1.
    The complement of a strongly regular graph of a (negative) Latin square type is a strongly regular graph of the same type.
  • 2.
    The union of two edge-disjoint and commuting strongly regular graphs of the same type is again a strongly regular graph of the same type.
Proof 2
The first claim is easily checked using the eigenvalues.
As for the second claim, let $A_{1}$ and $A_{2}$ be two such graphs with parameters $v_{i}=n^{2}$, $k_{i}=t_{i}(n-1)$, and restricted eigenvalues $a_{i}=-t_{i}$ and $b_{i}=(n-t_{i})$ for all $i\in\{1,2\}$. By Lemma 2.2 or the second part of Lemma 2.3, the restricted eigenvalues of $A_{1}+A_{2}$ are $a_{1}+b_{2}=a_{2}+b_{1}=n-(t_{1}+t_{2})$ and $a_{1}+a_{2}=-(t_{1}+t_{2})$. Note that the valency of $A_{1}+A_{2}$ is $k_{1}+k_{2}=(t_{1}+t_{2})(n-1)$. Therefore, $A_{1}+A_{2}$ is a strongly regular graph of the same type.
@@ 4 Schemes with many (negative) Latin square type relations
Proof 6
We note that if all but possibly one relation is of the same type, then it follows from Lemma 2.4 that the remaining relation, which is the complement of the union of all others must also be of the same type, and hence the scheme is amorphic [14]. We may thus assume that there is at least one relation of strictly Latin square type, say $A_{1}$, and at least one relation of strictly negative Latin square type, say $A_{2}$.
@@ 7.1 Strongly regular idempotents
Secondly, $vE_{1}\circ E_{1}=q^{0}_{11}E_{0}+q^{1}_{11}E_{1}+q^{2}_{11}E_{2}=q^{2}_{11}I+\frac{1}{v}(q^{0}_{11}-q^{2}_{11})J+(q^{1}_{11}-q^{2}_{11})E_{1}$. By multiplying this equation entrywise, once by $B_{1}$ and once by $B_{2}$, we obtain that $x^{2}-(q^{1}_{11}-q^{2}_{11})x-(q^{0}_{11}-q^{2}_{11})=0$ has solutions for $x=a,b$. Thus, $ab=q^{2}_{11}-q^{0}_{11}=q^{2}_{11}-m$ and $a+b=q^{1}_{11}-q^{2}_{11}$.